OpenStudy (anonymous):

write the product or quotient in standard form -8i(3-i)

OpenStudy (ybarrap):

8i(3-i)=24i-8 i*i=24i-8(-1)=8+24i

OpenStudy (anonymous):

so is the answer 8+24i ?

OpenStudy (ybarrap):

yes, do you understand why?

OpenStudy (anonymous):

no can you help explain?

OpenStudy (ybarrap):

do you know that i * i = -1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

o wait i get it! thanks!

OpenStudy (ybarrap):

i = sqrt(-1) so i * i = sqrt(-1) * sqrt(-1) = sqrt(-1)^2 =-1 I didn't see the negative in front of -8i(3-i). So take my answer above and negate it. Answer should be -8-24i -8 * i * 3 = -24i See that?